OpenStudy (anonymous):

PLEASE HELP!!! What is the simplified form of x/y^5 over x^2/y^2

OpenStudy (anonymous):

\[\frac{ x }{ y ^{5} } \over \frac{ x ^{2} }{ y ^{2} }\]

OpenStudy (anonymous):

you can rearrange that as\[\frac{ xy^2 }{ x^2y^5 }\] can you simplify that?

OpenStudy (anonymous):

a) x^3/x^7 b) 1/xy^3 c) xy^3 d) x/y^3

OpenStudy (psymon):

This is a regular division of fractions problem, just written in a funky manner. If you recall, when you divide fractions you flip the second fraction then multiply straight across. So this is (x/y^5) divided by (x^2/y^2). So I can rewrite that in the normal way without having the fractions stacked. Now by flipping the second fraction and multiplying, you have (x/y^5) * (y^2/x^2) which then combine to (xy^2/x^2y^5) From here you can simplify by regular division of exponents. one x on top and 2 x's on bottom leaves only 1 x on bottom. 2 y's on top and 5 y's on bottom leaves 3 y's on bottom. So that division simplifies to (1/xy^3)

OpenStudy (anonymous):

thank you!